# DetailMaster: Can Your Text-to-Image Model Handle Long Prompts?

### v1

@@ 7 Conclusion
Appendix A, Spatial partitioning scheme: We present our Spatial Partitioning Scheme for categorizing bounding boxes into positional regions and validate its superior performance over the conventional nine-grid partitioning method.
@@
Appendix B, Configurations of evaluated models: We detail the configurations of all evaluated models, including image resolutions and key inference hyperparameters.
@@
Appendix C, Time consumption and resource utilization: We document the computational costs associated with dataset construction and model evaluation, including detailed time consumption and GPU resource utilization.
Appendix D, Detailed human evaluation results: We present detailed human evaluation results, validating our benchmark’s high reliability.
@@
Appendix E, Evaluation pipeline details: We provide a comprehensive technical specification of our evaluation pipeline, systematically detailing the methodologies employed to assess accuracy across four critical dimensions: “Character Attributes”, “Character Locations”, “Scene Attributes”, and “Entity Relationships”.
@@
Appendix F, Randomness analysis for evaluation with LLMs: We present a randomness analysis for our LLM-based evaluation pipeline, empirically confirming the reproducibility and robustness of our evaluation protocol through random seed experiments.
Appendix G, Negative correlation between prompt length and attribute alignment: We present the correlation analysis between prompt length and attribute accuracy in generated characters, demonstrating a consistent negative relationship between prompt length and attribute alignment fidelity.
@@
Appendix H, Mini DetailMaster Benchmark: We introduces a mini-benchmark (800 detail-rich long prompts) that maintains evaluation consistency with the full benchmark while significantly reducing resource requirements, enabling rapid model assessment.
@@
Appendix I, Limitations: We examine the limitations of our work, specifically addressing constraints in both the characteristics of our evaluation data and evaluation pipeline.
@@
Appendix J, Case Study: We present representative case studies, and validate the accuracy of our comparative evaluations in the main text.
Appendix A Spatial partitioning scheme
@@
Our spatial partitioning scheme defines nine distinct regions within a normalized 1×1 coordinate space, where each region is represented as [$x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$] with ($x_{0}$, $y_{0}$) denoting the top-left coordinates and ($x_{1}$, $y_{1}$) the bottom-right coordinates. Specifically, we define: the upper part as [0.3, 0, 0.7, 0.5], the lower part as [0.3, 0.5, 0.7, 1], the left part as [0, 0.3, 0.5, 0.7], the right part as [0.5, 0.3, 1, 0.7], and the middle part as [0.3, 0.3, 0.7, 0.7]. Additionally, we establish four corner regions: the upper left [0, 0, 0.4, 0.4], lower left [0, 0.6, 0.4, 1], upper right [0.6, 0, 1, 0.4], and lower right [0.6, 0.6, 1, 1]. As illustrated in Figure 4, this partitioning scheme provides systematic coverage of the image space while maintaining clear semantic correspondence with spatial relationships. With the nine regions, we first normalize the character’s bounding box coordinates, then calculate the proportional area overlap between the normalized box and each of the nine spatial regions. The final position classification is determined by selecting the region that covers at least 75% of the character’s bounding box area and shows the highest proportional overlap among all qualifying regions. For cases where no region meets the 75% coverage threshold (indicating boundary straddling), we assign a “null” location designation.
@@
To validate the effectiveness of our nine-region partitioning scheme, we conduct a comparative analysis between the conventional 3×3 grid method and our proposed scheme by examining the proportion of ambiguous character positions (marked as “null”) at region boundaries. The experimental results show: while the traditional method classify 7,821 out of 11,816 characters (66.19%) as “null” due to boundary ambiguity, our partitioning scheme reduce this number to only 3,862 (32.68%). This reduction provides empirical evidence that our scheme effectively prevents characters from being situated near region boundaries and extracts more valid positional descriptions.
Additionally, we employ the MLLM Qwen2.5-VL-7B-Instruct to verify the consistency between our position descriptions and actual character locations, following the same method as our evaluation pipeline: we highlight characters with red bounding boxes and provide both the box coordinates and our position descriptions to the MLLM for verification. Comparative analysis reveal that while the conventional 3×3 grid method achieved 95.91% accuracy, our proposed method attain 95.52% accuracy. This marginal difference shows that our scheme successfully maintains comparable localization accuracy while significantly increasing the number of detectable character positions, confirming that our partitioning strategy achieves an optimal balance between precision and coverage.
Appendix B Configurations of evaluated models
In this section, we detail the configurations set for image generation across different models in our evaluation. Regarding the output resolution, we adopt 512×512 pixels for SD1.5 and its derivative models to align with their training settings, while setting 1024×1024 pixels for all other models. For inference steps (num_inference_steps), we follow each model’s reference implementations: 100 steps for DeepFloyd IF and ParaDiffusion, 70 steps for ELLA, and 50 steps for remaining models. The guidance scale parameters are similarly configured according to official recommendations: 11 for ELLA, 7 for DeepFloyd IF, 5 for SD-XL, 3.5 for both FLUX and SD3.5, and 7.5 for all other models.
@@ Appendix B Configurations of evaluated models
For LLM Blueprint, we employ Qwen2.5-7B-Instruct as the LLM to perform attribute extraction and position extraction from long prompts. For Deepfloyd, “IF-I-XL-v1.0” serves as the first-stage model, “IF-II-L-v1.0” serves as the second-stage model, and “stable-diffusion-x4-upscaler” serves as the upscaler.
@@
Appendix C Time consumption and resource utilization
@@
The evaluation pipeline typically consumes 10 hours using a single NVIDIA L20 GPU, with duration positively correlated with generated image quality.
@@
Regarding the GPU memory, both the first stage of our data construction pipeline and the evaluation pipeline require 20GB-39GB of GPU memory. When executing the second stage of our data construction pipeline across 8 GPUs, each GPU shows memory usage ranging from 26GB to 38GB. The variation in memory requirement arises due to the different values of the “cache_max_entry_count” parameter, ranging from 0.01 to 0.8. Additionally, the previously reported time consumption is measured with the “cache_max_entry_count” parameter set to 0.8.
Appendix D Detailed human evaluation results
@@
For the “Scene Attributes” (covering background, lighting, and style conditions), we involve three systematic checks: 1) Categorical Verification - validating that the annotated “scene_attribute_content” correctly describes the annotated visual condition; 2) Scene_Attribute-Image Alignment - determining whether the annotated “scene_attribute_content” aligns with the source image; and 3) Prompt Inclusion - determining whether the annotated “scene_attribute_content” is unambiguously specified in the final prompt. The results are shown in Table 7.
@@
Appendix E Evaluation pipeline details
We design a multi-stage evaluation pipeline to rigorously assess compositional text-to-image generation. The pipeline initiates by instructing target text-to-image models with prompts from DetailMaster to produce corresponding image sets, followed by automated main character detection and localization within the images. Subsequently, we perform sequential evaluations across four critical dimensions: 1) “Character Attributes”, 2) “Character Locations”, 3) “Scene Attributes” , and 4) “Entity Relationships”. This section provides comprehensive technical specifications of our evaluation pipeline, detailing the methodologies and validation mechanisms implemented for each evaluation dimensions.
Character presence identification.
Initially, we implement a dual-model verification system to identify which characters are successfully generated. We simultaneously employs an MLLM and the open-set detector YOLOE-11L for object detection. A character is considered successfully generated only when both models confirm its presence. For the generated characters, we further determine their locations to assist subsequent attribute evaluation. We calculate the IOU score between the two bounding boxes detected. If the score exceeds 0.7, we consider the boxes as overlapping and select the larger bounding box. Otherwise, we employ the vision-language model BLIP to compute image-text matching scores between both bounding box regions and the character name, selecting the box with the higher score. Identifying character presence can accelerate the whole evaluation process by eliminating redundant computations for failed generations. And it also avoids repetitive detection steps as the positional information can be reused.
“Character Attribute” evaluation.
For character attribute evaluation, we sequentially process each identified character and extract sub-images based on their positional information. Subsequently, we employ an MLLM to evaluate every attribute in the pre-built attribute list for each character. The MLLM provides binary judgments (“yes” or “no”) for each attribute based on the sub-image content. We then quantify the number of correctly generated attributes. Finally, the character attribute accuracy is calculated separately across three distinct categories: objects, animals, and persons, to enable category-specific performance analysis.
“Character Location” evaluation.
For character location evaluation, we sequentially process each identified character and mark its position in the generated image with a red bounding box. The bounding box information are then provided to the MLLM along with explicit instructions that the red box marks the target character. And the MLLM is prompted to determine whether the character precisely aligns with the position specified in its “Character Location” attribute, responding with a yes/no judgment. The final character location accuracy is calculated as the proportion of correctly positioned characters to the total number of characters who have “Character Location” attribute.
“Scene Attribute” evaluation & “Entity Relationship” evaluation.
For scene attribute evaluation, we simply ask the MLLM to determine whether the generated images satisfy each specified “Scene Attribute” with yes/no responses. The accuracy rates are calculated separately for background, lighting, and style attributes. Regarding entity relationship evaluation, for each generated image, we first mark all identified characters in the image with red bounding boxes. We then provide the MLLM with the marked image and the list of identified characters, and subsequently ask it to evaluate each attribute in the “Explicit Spatial/Interactive Relationships” checklist, determining whether the specified entity relationship aligns the image content.
Appendix F Randomness analysis for evaluation with LLMs

### v2

@@ 1. Core Methodology and Implementation Details:
  • Appendix B, Spatial Partitioning Scheme: We present our Spatial Partitioning Scheme for categorizing bounding boxes into positional regions and validate its superior performance over the conventional nine-grid partitioning method.
@@
  • Appendix C, Configurations of Evaluated Models: We detail the configurations of all evaluated models, including image resolutions and key inference hyperparameters.
@@
  • Appendix D, Time Consumption and Resource Utilization: We document the computational costs associated with dataset construction and model evaluation, including detailed time consumption and GPU resource utilization.
  • Appendix E, LLM Prompt Design for Our Pipeline: We detail the specific prompt designs employed throughout our LLM-driven pipeline. The effectiveness of our data construction and model evaluation stages relies heavily on these carefully crafted prompts, which are engineered to elicit precise and reliable outputs from the LLM.
@@
  • Appendix F, Evaluation Protocol: We provide a comprehensive technical specification of our evaluation pipeline, systematically detailing the methodologies employed to assess accuracy across four critical dimensions: “Character Attributes”, “Character Locations”, “Scene Attributes”, and “Entity Relationships”.
2. Robustness of Our Data Construction and Evaluation Protocol:
  • Appendix G, Detailed Human Evaluation Results: We present detailed human evaluation results, validating our benchmark’s high reliability.
  • Appendix H, Mitigating the Impact of Inherent LLM Biases via Auxiliary Techniques: We present an analysis of our pipeline against LLM-induced biases, empirically validating that our suite of auxiliary techniques ensures the objectivity and integrity of the data construction and evaluation protocol.
  • Appendix I, Robustness of the Single-MLLM Evaluation Protocol: We present a robustness analysis of our LLM-based evaluation pipeline by re-running all evaluations with an alternative evaluator (InternVL3-9B). The results empirically demonstrate the robustness of our protocol: while a different evaluator may introduce score inflation, the relative model rankings and overarching analytical trends are preserved.
@@ 2. Robustness of Our Data Construction and Evaluation Protocol:
  • Appendix J, Randomness Analysis for Evaluation with LLMs: We present a randomness analysis for our LLM-based evaluation pipeline, empirically confirming the reproducibility and robustness of our evaluation protocol through random seed experiments.
3. Supplementary Resources and Materials:
  • Appendix K, Compatibility of Our Framework with Other Evaluation Metrics: We demonstrate the compatibility of our benchmark by empirically evaluating its outputs with established external metrics, confirming its utility as a specialized component within broader, multi-faceted evaluation pipelines.
@@ 3. Supplementary Resources and Materials:
  • Appendix L, Mini DetailMaster Benchmark: We introduces a mini-benchmark (800 detail-rich long prompts) that maintains evaluation consistency with the full benchmark while significantly reducing resource requirements, enabling rapid model assessment.
4. Further Analysis and Insights:
  • Appendix M, Statistical Analysis of Metric Difficulty: we present a systematic analysis of metric difficulty, quantitatively establishing a clear hierarchy from Character Locations (most difficult) to Scene Attributes (easiest). Furthermore, our qualitative investigation of failure cases reveals common error patterns that illuminate the underlying weaknesses of the models.
  • Appendix N, Negative Correlation between Prompt Length and Attribute Alignment: We present the correlation analysis between prompt length and attribute accuracy in generated characters, demonstrating a consistent negative relationship between prompt length and attribute alignment fidelity.
@@ 4. Further Analysis and Insights:
  • Appendix O, Limitations: We examine the limitations of our work, specifically addressing constraints in both the characteristics of our evaluation data and evaluation pipeline.
@@
  • Appendix P, Case Studies: We present representative case studies, and validate the accuracy of our comparative evaluations in the main text.
Appendix B Spatial Partitioning Scheme
@@
Our spatial partitioning scheme defines nine distinct regions within a normalized 1×1 coordinate space, where each region is represented as $x_{0}$, $y_{0}$, $x_{1}$, $y_{1}$] with ($x_{0}$, $y_{0}$) denoting the top-left coordinates and ($x_{1}$, $y_{1}$) denoting the bottom-right coordinates. Specifically, we define: the upper part as [0.3, 0, 0.7, 0.5], the lower part as [0.3, 0.5, 0.7, 1], the left part as [0, 0.3, 0.5, 0.7], the right part as [0.5, 0.3, 1, 0.7], and the middle part as [0.3, 0.3, 0.7, 0.7]. Additionally, we establish four corner regions: the upper left [0, 0, 0.4, 0.4], lower left [0, 0.6, 0.4, 1], upper right [0.6, 0, 1, 0.4], and lower right [0.6, 0.6, 1, 1]. As illustrated in Figure 4, this partitioning scheme provides systematic coverage of the image space while maintaining clear semantic correspondence with spatial relationships. With the nine regions, we first normalize the character’s bounding box coordinates, then calculate the proportional area overlap between the normalized box and each of the nine spatial regions. The final position classification is determined by selecting the region that covers at least 75% of the character’s bounding box area and shows the highest proportional overlap among all qualifying regions. For cases where no region meets the 75% coverage threshold (indicating boundary straddling), we assign a “null” location designation.
@@
To validate the effectiveness of our partitioning scheme, we conduct a comparative analysis between the conventional 3×3 grid method and our proposed scheme by examining the proportion of ambiguous character positions (marked as “null”) at region boundaries. The experimental results show: while the traditional method classify 7,821 out of 11,816 characters (66.19%) as “null” due to boundary ambiguity, our partitioning scheme reduce this number to only 3,862 (32.68%). This reduction provides empirical evidence that our scheme effectively prevents characters from being situated near region boundaries and extracts more valid positional descriptions.
Additionally, we employ the MLLM Qwen2.5-VL-7B-Instruct to verify the consistency between our position descriptions and actual character locations, following the same method as our evaluation pipeline: we highlight characters with red bounding boxes and provide both the box coordinates and the position descriptions to the MLLM for verification. Comparative analysis reveal that while the conventional 3×3 grid method achieved 95.91% accuracy, our proposed method attain 95.52% accuracy. This marginal difference shows that our scheme successfully maintains comparable localization accuracy while significantly increasing the number of detectable character positions, confirming that our partitioning strategy achieves an optimal balance between precision and coverage.
Appendix C Configurations of Evaluated Models
In this section, we detail the configurations for image generation across different models in our evaluation. Regarding the output resolution, we adopt 512×512 pixels for SD1.5 and its derivative models to align with their training settings, while setting 1024×1024 pixels for all other models. For inference steps (num_inference_steps), we follow each model’s reference implementations: 100 steps for DeepFloyd IF and ParaDiffusion, 70 steps for ELLA, and 50 steps for remaining models. The guidance scale parameters are similarly configured according to official recommendations: 11 for ELLA, 7 for DeepFloyd IF, 5 for SD-XL, 3.5 for both FLUX and SD3.5, and 7.5 for all other models.
@@ Appendix C Configurations of Evaluated Models
For LLM Blueprint, we employ Qwen2.5-7B-Instruct as the LLM to perform attribute extraction and position extraction from long prompts. For Deepfloyd IF, “IF-I-XL-v1.0” serves as the first-stage model, “IF-II-L-v1.0” serves as the second-stage model, and “stable-diffusion-x4-upscaler” serves as the upscaler.
@@
Appendix D Time Consumption and Resource Utilization
@@
The evaluation pipeline typically consumes 10 hours using a single NVIDIA L20 GPU, with duration positively correlated with the quality of generated images.
@@
Regarding the GPU memory, both our data construction pipeline’s first stage and the evaluation pipeline require 20GB-39GB of GPU memory. When executing the second stage of our data construction pipeline across 8 GPUs, each GPU shows memory usage ranging from 26GB to 38GB. The variation in memory requirement arises due to the different values of the “cache_max_entry_count” parameter, ranging from 0.01 to 0.8. Additionally, the previously reported time consumption is measured with the “cache_max_entry_count” parameter set to 0.8.
Appendix E LLM Prompt Design for Our Pipeline
In this section, we present the prompt designs for the LLMs (MLLMs) used in our data construction and model evaluation phases.
In the “Main Character Identification” stage of our data construction process, we employ a few-shot approach to construct prompts that guide the model to respond in a JSON format. The prompt informs the model that it will be provided with “an image and its corresponding description”, and its task is to “identify and count the main characters in the image”. During prompt tuning, we experimented with zero-shot prompting but observed unsatisfactory performance. We also tried omitting the character count, which similarly degraded results.
For the “Character Localization” stage, we use the prompt “Please only provide the bounding box coordinates of the region CHARACTER describes”, where CHARACTER refers to the identified main character from the previous stage. While Qwen2.5-VL-Instruct generally adheres to this instruction, occasional localization errors lead us to deploy an open-set detector YOLOE-11L to vote for the correct bounding box coordinates.
When extracting character attributes, we provide the model with the relevant image captions and, when necessary, highlight or crop the target regions of the characters while also supplying examples of the attributes to be extracted. During prompt tuning, we evaluated two alternative approaches: one eliminating image captions and another removing cropping/highlighting operations. Both configurations resulted in fewer extracted attributes.
In the “Prompt Refinement” stage, we iteratively refine the image captions by prompting the model to supplement missing attributes: “Some of the CATEGORY attribute information in the image description is missing. Your task is to supplement the missing CATEGORY attribute information into the image description.” Here, CATEGORY denotes different attribute types. For the “Filtering” stage, we prompt the model to determine which attributes are absent from the caption: “Your task is to determine whether the given image description includes the description of this particular CATEGORY feature of the main character.”
For evaluating “Character Attribute”, we prompt the MLLM to analyze each attribute individually, using cropped images of the characters: “Please analyze the main character in this image, specifically the CHARACTER. Please determine whether ATTRIBUTE is one of its characteristics or is associated with it.” Here, CHARACTER and ATTRIBUTE are the specific character name and the attribute being evaluated. In the “Character Location” evaluation, we highlight the image and provide positional coordinates, prompting the MLLM to assess whether the object’s location matches the reference. Regarding the feature evaluation for the entire image, the MLLM directly judges based on the image content and attribute categories.
Appendix F Evaluation Protocol
We design a multi-stage evaluation pipeline to rigorously assess compositional text-to-image generation. The pipeline initiates by instructing text-to-image models with prompts from DetailMaster to produce corresponding image sets, followed by automated main character detection and localization within the images. Subsequently, we perform sequential evaluations across four critical dimensions: 1) “Character Attributes”, 2) “Character Locations”, 3) “Scene Attributes” , and 4) “Entity Relationships”. Regarding the evaluator, we employ Qwen2.5-VL-7B-Instruct as the MLLM throughout the evaluation process.
This section provides comprehensive technical specifications of our evaluation pipeline, detailing the methodologies and validation mechanisms implemented for each evaluation dimensions.
Character presence identification.
Initially, we implement a dual-model verification system for character detection, where a character is deemed successfully generated only if its presence is confirmed by both an MLLM and the YOLOE-11L open-set detector. For the confirmed characters, we also record their box coordinates to assist the subsequent attribute evaluation. Regarding how to determine the box coordinates, we calculate the IoU score between the two bounding boxes detected. If the score exceeds 0.7, we consider the boxes as overlapping and select the larger bounding box. Otherwise, we employ the vision-language model BLIP to compute image-text matching scores between both bounding box regions and the character name, selecting the box with a higher score. Identifying character presence can accelerate the whole evaluation process by eliminating redundant computations for failed generations. And it also avoids repetitive detection steps as the positional information can be reused.
“Character Attribute” evaluation.
Our methodology for character attribute evaluation involves three main steps. First, we crop sub-images of individual present characters using their positional information. Second, each sub-image is processed by an MLLM, which renders a binary (“yes” or “no”) judgment on the presence of each specified attribute. Third, we measure performance by calculating the accuracy of attribute generation. This metric is computed separately for objects, animals, and persons to allow for a detailed, category-level performance analysis.
“Character Location” evaluation.
For character location evaluation, we sequentially process each identified character and mark its position in the generated image with a red bounding box. The bounding box information are then provided to the MLLM along with explicit instructions that the red box marks the target character. And the MLLM is prompted to determine whether the character location precisely aligns with the position description, responding with a yes/no judgment. The final character location accuracy is calculated as the proportion of correctly positioned characters to the total number of characters who have “Character Location” attribute.
“Scene Attribute” evaluation & “Entity Relationship” evaluation.
For scene attribute evaluation, we ask the MLLM to determine whether the generated images satisfy each specified “Scene Attribute” with yes/no responses. The accuracy rates are calculated separately for background, lighting, and style attributes. Regarding entity relationship evaluation, for each generated image, we first mark all identified characters in the image with red bounding boxes. We then provide the MLLM with the marked image and the list of identified characters, and subsequently ask it to evaluate each attribute in the “Spatial/Interactive Relationships” checklist, determining whether the specified entity relationship aligns with the image content.
Appendix G Detailed Human Evaluation Results
@@
For the “Scene Attributes” (covering background, lighting, and style conditions), we involve three systematic checks: 1) Categorical Verification - validating that the annotated “scene_attribute_content” correctly describes the annotated visual condition (i.e., background, lighting, and style); 2) Scene_Attribute-Image Alignment - determining whether the annotated “scene_attribute_content” aligns with the source image; and 3) Prompt Inclusion - determining whether the annotated “scene_attribute_content” is unambiguously specified in the final prompt. The results are shown in Table 7.
@@
Appendix H Mitigating the Impact of Inherent LLM Biases via Auxiliary Techniques
Our pipeline leverages advanced LLMs for data construction and model evaluation. A critical consideration, therefore, is the potential influence of inherent biases from these models. In this section, we address this concern and present supplementary experiments to demonstrate that our pipeline is robust against such biases. We argue that the extensive integration of auxiliary techniques at each stage effectively mitigates these potential influences, ensuring the quality and objectivity of our dataset and evaluation process.
H.1 On Potential Bias in LLM-based Data Construction
A primary concern is that sole reliance on LLMs for data generation could introduce inherent biases of the models. To counteract this, we integrate a suite of auxiliary techniques into our data construction workflow, including open-set object detection, image cropping, image highlighting, and few-shot prompting. These techniques serve to ground the MLLM’s outputs in objective image features rather than learned priors, enhancing the accuracy and diversity of the generated data. This multi-faceted approach ensures that our dataset is not exclusively dependent on the generative capabilities of any single LLM, thereby insulating it from model-specific biases.
Furthermore, our selection of MLLM is also a deliberate step in quality control. We benchmark three leading models—Qwen2.5-VL-Instruct, InternVL2.5 (Chen et al., 2024), and LLaVA-OneVision (Li et al., 2024b)—and observe that Qwen2.5-VL-Instruct exhibits more rigorous logical reasoning and generates more precise outputs. Consequently, we select it as the primary MLLM for our data construction pipeline.
H.2 On Potential Bias in LLM-based Evaluation
A second concern is a potential evaluation bias favoring text-to-image models that use LLM-based text encoders, given that our data is constructed with LLMs. This concern is unfounded for two main reasons. First, the LLMs employed as text encoders in the evaluated models (e.g., Llama-2, T5-XL) are architecturally distinct and typically older than the LLM chosen for our data construction, Qwen2.5-VL-Instruct. (For the LLM Blueprint model specifically, while we employ Qwen2.5-7B-Instruct for object name extraction from prompts, this LLM functions only as a pre-processing component and is not the text encoder used during image generation.)
Second, as previously detailed, our extensive use of auxiliary techniques ensures the rich diversity and factual accuracy of our DetailMaster dataset. This effectively decouples the dataset from an over-reliance on the specific characteristics of Qwen2.5-VL-Instruct, minimizing the potential inherent biases that could favor a particular model architecture during evaluation.
H.3 Experimental Validation
To empirically validate our claims, we conduct additional ablation studies analyzing the impact of our auxiliary techniques and the choice of LLMs on the data construction process.
Impact of auxiliary techniques on data quality.
We first examine the effect of the deployed open-set object detector during the Character Localization stage. Our findings indicate that this step corrects 16.6% of bounding box coordinates proposed by the MLLM, enhancing localization accuracy.
Next, we investigate the consequences of omitting image cropping during the Character Attribute Extraction stage. Specifically, we employ Gemini-2.5-Flash (Comanici et al., 2025) to evaluate the accuracy of each generated annotation. While the accuracy remains relatively stable, we observe a 15.1% reduction in the number of extracted attributes. This demonstrates that our image cropping is crucial for enabling the MLLM to perform a more comprehensive and detailed analysis.
Impact of MLLM choice on data generation.
As previously mentioned, Qwen2.5-VL-Instruct shows relatively strict inference behavior, which our new experiment further confirms. We replicate the Character Attribute Extraction pipeline using InternVL3-9B (Zhu et al., 2025) and evaluate the resulting annotations with Gemini-2.5-Flash. The results indicate a 3.4% decrease in accuracy. Moreover, due to format errors in InternVL3-9B’s outputs (e.g., failure to adhere to list format), the quantity of generated attributes decreases by 24.5%.
In conclusion, this section has rigorously addressed the critical issue of LLM-induced bias within our pipeline. Through a series of targeted experiments, we have demonstrated conclusively that our suite of auxiliary techniques is indispensable for achieving high data quality. These methods not only rectify potential inaccuracies but also enrich the detail of our generated data. The performance degradation observed when substituting our carefully selected MLLM further underscores the efficacy of our overall approach. Therefore, we assert with confidence that our pipeline is robust against LLM-specific biases, establishing the DetailMaster dataset as a reliable and impartial resource for the community.
Appendix I Robustness of the Single-MLLM Evaluation Protocol
It is a common practice in existing benchmarks for text-to-image models to employ a single MLLM for evaluation. For instance, T2I-CompBench utilizes only Minigpt-4 (Zhu et al., 2023) for feature verification, while DPG-Bench relies on a single mPLUG-large model (Li et al., 2022a) for characteristic assessment. There are two primary reasons why benchmark works typically deploy only a single MLLM: (1) Deploying multiple MLLMs consumes significant time and GPU computing resource; (2) Adapting prompts, auxiliary tools, or evaluation frameworks for different MLLMs would lead to overly cumbersome inference code.
Actually, our selection of Qwen2.5-VL-7B-Instruct is based on comparisons. We evaluate it against InternVL2.5 (Chen et al., 2024) and LLaVA-OneVision(Li et al., 2024b), and find that Qwen2.5-VL-7B-Instruct demonstrates stricter judgment criteria. It shows superior accuracy in identifying object features while exhibiting strong adherence to prescribed output formats. In contrast, both InternVL2.5 and LLaVA-OneVision exhibit a tendency to respond “yes” more readily during feature evaluation, which consequently leads to inflated scores.
We conduct a new set of experiments, re-running our entire evaluation pipeline using InternVL3-9B. The results are presented in Table 9.
Table 9: Evaluation results on the DetailMaster Benchmark (with InternVL3-9B).
Model Character Attributes Character Locations Scene Attributes Entity Relationships
Object Animal Person Background Light Style
SD1.5 22.62 24.03 17.84 11.90 33.48 77.20 83.90 12.99
SD-XL 36.41 44.03 29.19 20.97 38.62 79.82 69.65 19.63
DeepFloyd IF 54.75 57.92 45.25 33.11 37.12 79.66 85.18 28.30
SD3.5 Large 67.57 78.99 55.38 59.26 95.66 95.11 96.28 67.18
FLUX.1-dev 71.42 79.42 55.61 67.13 98.85 98.78 96.01 74.41
Gemini 2.0 Flash 72.37 80.08 54.65 69.31 98.72 98.62 97.51 77.27
GPT Image-1 73.44 83.81 59.02 69.95 99.51 99.11 95.75 82.51
LLM4GEN 24.25 30.76 21.13 14.66 35.82 77.86 55.06 16.23
LLM Blueprint 22.02 22.90 18.65 28.26 66.03 87.79 65.30 20.61
ELLA 41.92 49.76 34.93 27.99 57.82 84.44 42.66 26.74
LongAlign 40.54 47.52 26.53 25.98 86.26 93.00 84.58 31.61
ParaDiffusion 51.68 52.84 39.11 45.88 90.45 95.45 68.13 56.88
As shown in the table, with InternVL3-9B, all models show varying degrees of score improvement. This observation aligns with our previous experimental findings on InternVL2.5, suggesting that using InternVL3-9B for evaluation is not sufficiently rigorous, as it tends to favor “yes” responses when assessing ambiguous features.
However, despite the overall score improvement, the comparative rankings among the models and the general score trends remain unchanged. For general-purpose T2I models, more advanced models consistently achieve higher scores. Similarly, long-prompt optimized T2I models outperform the baseline models SD1.5 and SD-XL.
Regarding additional details, for instance:
  • Top-tier proprietary models (GPT Image-1, Gemini 2.0 Flash) still outperform all other models, with more advanced models consistently achieving higher scores.
  • Long-prompt-specific models (ParaDiffusion, LongAlign, ELLA) still show clear advantages over baselines like SD-XL on complex attributes.
  • ParaDiffusion continues to excel in Character Locations and Entity Relationships compared to other long-prompt-specific models, validating the effectiveness of task-specific adaptive training (Jiao et al., 2024; Xu et al., 2025).
  • We also compute the Kendall’s Tau Correlation Coefficient between the evaluation results obtained with the two MLLMs. The correlation coefficients for each metric are as follows: 0.879, 0.909, 0.909, 0.909, 0.970, 1.000, 0.879, and 0.970. These consistently high values (all approaching 1.0) reflect a strong positive correlation between the two MLLM evaluators, confirming the consistent scoring trends across the evaluated models.
This cross-evaluator consistency provides strong evidence that our benchmark’s conclusions are robust and not an artifact of the specific MLLM used.
Appendix J Randomness Analysis for Evaluation with LLMs